OpenStudy (anonymous):

Medal and Fan Find the GCF of the terms of the polynomial. 16x3 + 28x5y A. 8x5 B. 112x5y C. 4x5y D. 4x3

OpenStudy (anonymous):

GCF = Greatest Common Factor, so first you want to factor out the two terms in the polynomial

OpenStudy (anonymous):

\[What\]

OpenStudy (anonymous):

do you know how to factor?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

So for the first one, you would have 16x^3 = 2*2*2*2*x*x*x

OpenStudy (anonymous):

can you do the other?

OpenStudy (anonymous):

No It's Hard

OpenStudy (anonymous):

I'm Completly LOST

OpenStudy (anonymous):

28x^5y = 2* 2 * 7*x*x*x*x*x*y

OpenStudy (anonymous):

OK, that is how you factor what you want to do after that is look to see where the greatest factors are

OpenStudy (anonymous):

We see that there are 2 2's in both of them and 3 x's

OpenStudy (anonymous):

Is it B

OpenStudy (anonymous):

no, which answer would it be?

OpenStudy (anonymous):

A

OpenStudy (anonymous):

no, 2*2=4 and x*x*x = x^3 so your answer is 4x^3

OpenStudy (anonymous):

does that help?

OpenStudy (anonymous):

D?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thanks Man

OpenStudy (anonymous):

Glad to help

OpenStudy (anonymous):

I have more :P

OpenStudy (anonymous):

Don't forget to close the question